OpenStudy (anonymous):

How do you find an exact number in Trig? I have problems such as finding sin45 when it gives me a triangle with 2 side lengths and two angles

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

there are 2right angled triangles that are worth remembering 1.is an isioceles triangle with two sides about the right angle both = 1 and the hypotenuse(lonfest side) = Sqrq 2 The two equal angles = 45 degrees. So if u sketch this triabgle u will see that tan 45 = 1, cos and sin 45 = / sqrt2 The other has angles 30 and 60 and 90 degrees. Here the longest side = 2, shortest - 1 and other side = sqrt 3 so if you sketch these u will se that cos 60, = 1/2, sin 30 = 1/2, cos 30 = sqrt3 / 2 and sin 60 = sqrt3 / 2. These are exact values

OpenStudy (anonymous):

cos and sin 45 = 1 / sqrt2 - this also is an exact vale

OpenStudy (anonymous):

any moe questions feel free to ask

OpenStudy (anonymous):

So how is it that when I do these problems I get the same answers for each type every time? Is that simply how they work out?

OpenStudy (anonymous):

well yes the 45 degree triangle is always the same shape whatever the lengths of the sides if the two perpenducular sides were bot 10 long the tan of 45 = 10/10 = 1

OpenStudy (anonymous):

in this case the hypotenuse will be 10*sqrt 2 and the sine 45 = 10/10*sqrt2 = 1/sqrt2

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

sorry, I was at school. I did get it figured out though :) thank you for your help!